OpenStudy (anonymous):

Please solve 8x − 2(x + 1) = 7x + 8

OpenStudy (anonymous):

8x-2x-2=7x+8 6x-2-7x-8=0 -x=10 x=10

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

you're welcome :)

OpenStudy (anonymous):

sorry, x = -10